OpenStudy (anonymous):

How do I solve this by completing the square: (x^2) - 7x - 3 = 0

OpenStudy (anonymous):

cube = 61

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

is it an A B C D question

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

OK. So to complete the square, you're going to try to make something of the form (x-a)^2, and then add a little to fix it.

OpenStudy (anonymous):

Yes, I did that with previous questions but when I tried it for this one, I got stuck.

OpenStudy (anonymous):



OpenStudy (merchandize):

x^2-7x-3=0 x^2-3x-4x-3=0 x(x-3)-4(x-3)=0 (x-3) (x-4) x=3,4 answer

OpenStudy (anonymous):

The final answer is supposed to be in surd form though